OpenStudy (anonymous):

A company rented a copy machine for $225 per month for 3 months. The company also paid $0.012 for each copy made. If the total cost of using the copy machine during this 3-month period was $1,107, how many copies did the company make? A 148,500 B 92,250 C 73,500 D 36,000

OpenStudy (anonymous):

This easy

OpenStudy (anonymous):

(225*3)+(0.012*C)= $1.107 Can u solve now?

OpenStudy (anonymous):

is D

OpenStudy (anonymous):

it's true . Thank u so much

OpenStudy (anonymous):

675+0.012C=1.107 C = (1.107-675)/0.012 C= 36.000